# Conformalized Bayesian Inference, with Applications to Random Partition Models

### v1

@@ 3.2 Credible region
with $\mathbb{P}_{\bm{\theta}^{2}\overset{\textnormal{iid}}{\sim}\Pi}$-probability at least $1-\delta$, implying $O(N^{-1/2})$ concentration around the desired coverage level $1-\alpha$. For a high-probability guarantee that the coverage merely exceeds $1-\alpha$, one may instead use the results in the references cited above.
@@
Before proceeding with illustrations of the conformal credible region procedure, we provide theoretical justification for the iid assumption on the calibration scores. Recall that we collect N:=TSN:=T-S calibration samples, and for simplicity re-index them as $\theta_{0},\theta_{M},\ldots,\theta_{M(N-1)}$, where $M$ acts as a thinning parameter determining which samples are included in the calibration set (that is, every $M$th sample from the sample is retained). While the iid assumption is crucial for obtaining a valid credible region under conformal prediction principles, in practice we often have access not to iid samples from $\Pi$, but to Markov-dependent samples produced by an MCMC sampler. However, such a sampler is typically designed to be $\Pi$-mixing, meaning that $\lim_{t\to\infty}\sup_{\theta_{0}\in\Theta}\|\mathcal{L}_{t}(\cdot\mid\theta_{0})-\Pi\|_{TV}=0,$ where $\mathcal{L}_{t}(\cdot\mid\theta_{0})$ denotes the distribution of $\theta_{t}$ conditional on the chain starting at $\theta_{0}$, and $\|\cdot\|_{TV}$ denotes the total variation (TV) distance between probability measures. Recall that, for two probability measures $\mu$ and $\nu$ on a sample space $\Omega$, the TV distance is defined as
@@
Under stronger assumptions, one can sometimes establish quantitative convergence, i.e., that the chain is $\Pi$-mixing at rate $\varepsilon_{t}$, meaning that $\sup_{\theta_{0}\in\Theta}\|\mathcal{L}_{t}(\cdot\mid\theta_{0})-\Pi\|_{TV}\leq\varepsilon_{t},$ for some decreasing sequence $(\varepsilon_{t})_{t\in\mathbb{N}}$; see Levin and Peres, (2017) for a standard and comprehensive treatment.
The following result shows that, as long as the chain is mixing, the joint law of the calibration scores $(s_{0},s_{M},\ldots,s_{M(N-1)}),$ provided that they are computed from a suitably thinned calibration sample $(\theta_{0},\theta_{M},\ldots,\theta_{M(N-1)}),$ approaches that of a posterior iid vector in TV distance (as $M\to\infty$).
@@
In practice, Proposition 2 ensures that, even if the sampled parameters are Markov dependent, [^5] Notice that, in our random partition application, the MCMC algorithm operates on a higher-dimensional space than $\Theta$ (i.e., including mixture parameters in addition to partitions). The resulting chain is then Markov on this larger space, but not necessarily when projected on $\Theta$. This does not affect Proposition 2, since the scores can be viewed as functions from the higher-dimensional state space to the real line. For simplicity, however, we state the result assuming the chain is defined directly on $\Theta$. appropriate thinning allows one to treat them as effectively iid. In fact, if the chain is $\Pi$-mixing at rate $\varepsilon_{t}$, one can even choose the spacing $M$ to achieve a desired coverage level in light of Proposition 1: by the definition of the TV distance in terms of bounded test functions, Proposition 2 implies that selecting $M$ large enough ensures an expected value of $\Pi(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2}))$ close to $1-\alpha$ (as $\bm{\theta}^{2}\mapsto\Pi(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2}))$ is bounded by 1). Similarly, using the set-based definition of TV distance, one can tune $M$ to obtain an approximate high-probability coverage statement.
We conclude this discussion by noting that the linear dependence on $N$ of the bound in Equation (7) arises from a rather crude argument that treats the calibration samples as ordered and Markov dependent. In practice, the credible set construction does not depend on order, so the samples may be randomly shuffled and regarded as exchangeable without loss of generality. From this viewpoint, a larger $N$ can even improve the TV bound, since shuffling more samples tends to reduce dependence more effectively. We do not pursue this refinement further, as it adds little methodological insight beyond the main point that, after appropriate thinning and burn-in, the calibration samples can be safely regarded as iid draws from the target posterior distribution.
@@
Consider the simulated data and MCMC output analyzed in Example 1. Based on the $\mathcal{D}_{VI}$–KDE calibration scores, we construct the credible set $\mathcal{C}_{0.9}(\bm{\theta}^{2})$. To assess its properties, we test whether the following partitions belong to it: (i) the “true” data partition shown in Figure 1(a); (ii) the same partition with the two leftmost clusters merged; and (iii) 1,000 randomly generated partitions, each drawn independently by first sampling $K$ from the empirical posterior distribution of the number of clusters (cf. Figure 1(b)), and then assigning each observation to one of $K$ clusters uniformly at random.
@@
We find that both partitions in (i) and (ii) are contained in the set. Given the point estimate shown in Figure 1(c), the inclusion of the collapsed true partition is unsurprising, as it closely resembles that estimate. More interestingly, the true partition itself is also included in the set, even though point estimation alone would not have led us to expect this, since the true and collapsed partitions are quite different. This indicates the presence of posterior uncertainty, likely in the form of multimodality, which we will examine further in the next subsection. Finally, none of the 1,000 randomly generated partitions fall within the credible region. This supports the view that the credible set does not achieve high coverage simply by encompassing many arbitrary partitions (e.g., those consistent with the marginal posterior on $K$ but otherwise unstructured), but rather by concentrating on regions of high posterior density. In other words, the $\mathcal{D}_{VI}$–KDE score yields a credible region that is both practically informative and well aligned with the posterior distribution on the space of partitions.
3.3 Multimodality analysis
@@
Having defined our approach to point estimation and credible region construction, we now turn to the last piece of the CBI program, that is to address multimodality in the posterior distribution over $\Theta$. Importantly, because the parameter spaces embodied by $\Theta$ are generally non-Euclidean, we require a notion of multimodality that is operationally meaningful within these structures. In particular, we have defined a density-like function, based on the training data, that can be evaluated at any $\theta\in\Theta$, and we can compute a distance $\mathcal{D}(\theta,\theta^{\prime})$ between any two parameters $\theta,\theta^{\prime}\in\Theta$. These two ingredients are sufficient to define a notion of multimodality grounded in the idea of density-based clustering. The latter provides a general framework, encompassing many algorithmic instantiations (Ester et al.,, 1996; Kriegel et al.,, 2011; Rodriguez and Laio,, 2014), based on the following principle: given a density function on $\Theta$ and a measure of discrepancy between its elements, one partitions the data into clusters such that (a) each cluster lies in a high-density region whose elements are mutually close, and (b) points from different clusters are far apart. A distribution is then said to be multimodal if random samples drawn from it fall into more than one cluster.
@@
Among the many algorithms following this idea, we take inspiration from the Density Peak Clustering (DPC) procedure by Rodriguez and Laio, (2014), which is simple, computationally efficient, and well suited to our setting. In particular, since the construction of the credible region already requires the NN $\mathcal{D}$-KDE scores for the calibration set, this algorithm can be applied with minimal additional computation. The only remaining step is to compute the pairwise distances $\mathcal{D}(\theta,\theta^{\prime})$ between all distinct calibration samples $\theta,\theta^{\prime}$; this has $O(N^{2})$ cost, but is parallelizable across each distance computation. Then, our procedure, which we term KDE-DPC, proceeds as follows:
@@
    For each calibration parameter $\theta\in\bm{\theta}^{2}$, compute
    $\delta(\theta):=\min_{\theta^{\prime}\in\bm{\theta}^{2}\,:\,s(\theta^{\prime};\bm{\theta}^{1})>s(\theta;\bm{\theta}^{1})}\mathcal{D}(\theta,\theta^{\prime}),$
    that is, the minimum distance from $\theta$ to any other parameter with higher density. For the parameter(s) with highest density (corresponding to our point estimator $\theta_{\star}$), set
    $\delta(\theta_{\star}):=\max_{\theta^{\prime}\in\bm{\theta}^{2}}\mathcal{D}(\theta_{\star},\theta^{\prime}).$
@@
    Identify the potential cluster centers as those calibration parameters exhibiting simultaneously high density $s(\theta;\bm{\theta}^{1})$ and large separation $\delta(\theta)$.
@@
We note that, if one were to follow Rodriguez and Laio, (2014) exactly, the density for each calibration sample $\theta$ would instead be estimated as the proportion of samples within a fixed-radius $\mathcal{D}$-neighborhood of $\theta$. While feasible, this would waste computation, since the KDE scores have already been computed, and it would ignore the information contributed by the training samples $\bm{\theta}^{1}$. Nevertheless, the overall approach is retained, as the $\mathcal{D}$-KDE remains distance-based, with the exponential kernel acting as a smoother version of a strict neighborhood.
The KDE-DPC procedure exhibits several appealing features. First, as illustrated in Example 3, the selection of representative parameter values corresponding to posterior modes reduces to a simple visual task: inspecting a decision graph where, for each calibration sample $\theta$, the quantities $\delta(\theta)$ (on the vertical axis) and $s(\theta;\bm{\theta}^{1})$ (on the horizontal axis) are plotted. The posterior modes are then identified as those samples that stand out as having abnormally large values of both quantities—that is, points lying in the north-east corner of the decision graph. This approach has the practical advantage of not requiring a pre-specified the number of modes, which is instead discovered together with the modes themselves. Second, the posterior weight associated with each mode can be naturally estimated as the proportion of calibration samples assigned to its cluster. Third, the method provides a straightforward way to classify calibration samples as inliers or outliers: samples lying in the north-west corner of the decision graph can be interpreted as outliers, since they are far from higher-density samples yet have low density themselves, while those located in the middle and south-east regions are inliers, as they exhibit moderate to high density and are close to representative modes. These considerations significantly enhance one’s understanding of the uncertainty encoded by the posterior distribution.
@@
We continue our analysis of the MCMC samples of random partitions introduced in Examples 1 and 2. Figure 2(a) displays the KDE-DPC decision graph, plotting for each calibration partition its separation $\delta$ and density $s(\cdot;\bm{\theta}^{1})$ values, with points colored by the number of clusters in the corresponding partition. [^6] A large number of two-cluster partitions in the calibration set coincide exactly with the partition corresponding to the north-east-most point, which explains the limited number of visible blue points. Two points (highlighted in red) clearly stand out in the north-east corner of the graph, identifying the representative modes. The corresponding partitions, shown in Figure 2(b) (ordered by density from left to right), indicate that the posterior concentrates substantial mass around two distinct configurations: one with two clusters, obtained by merging the two leftmost clusters of the “true” data partition, and another nearly identical to the “true” partition itself. [^7] See the Supplementary Material for an additional experiment using the same simulated data and a similar prior, where the posterior over partitions is instead found to be unimodal.
Interestingly, coloring by the number of clusters reveals that, while the posterior assigns non-negligible mass to partitions with four clusters, none of these emerge as representative modes. Instead, they appear either as inliers close to one of the two main modes (the yellow points at the center of the decision graph) or as outliers lying in low-density regions far from the modes (the yellow points in the north-west area of the decision graph). In this sense, the multimodality analysis suggests that, although the four-cluster configurations receive some posterior support under the PY mixture model used to fit the data, they are likely spurious from the standpoint of a parsimonious description of the uncertainty encoded in the posterior distribution.
@@
To conclude, we note that the output of our multimodality analysis based on the KDE-DPC procedure is analogous to that of Balocchi and Wade, (2025), who summarize the posterior over partitions using a Wasserstein-distance–minimizing mixture of point masses at representative partitions. In both approaches, the result is a set of posterior modes capturing meaningfully separated regions of posterior concentration in the partition space. A few advantages of our approach are worth highlighting. First, it only requires density scores (already computed for the credible region) and pairwise VI distances between calibration samples (also required by Balocchi and Wade, (2025)), both of which can be parallelized to achieve substantial computational gains. In contrast, the $k$-medoids-like algorithm of Balocchi and Wade, (2025) is inherently sequential. Second, their method requires specifying the number of modes in advance. To select a good value of this important hyperparameter, the authors propose running the algorithm for different choices and picking the one with the largest improvement in posterior approximation. Clearly, this approach may be computationally cumbersome and prone to overfitting to the available MCMC samples. Our KDE-DPC procedure, by contrast, automatically reveals the number of modes through visual inspection of the decision graph. Finally, Balocchi and Wade, (2025) propose several insightful techniques to characterize posterior uncertainty once the modes are identified, such as plotting the meet between any two representative partitions or examining the VI contribution of each data point. All of these analyses can naturally be performed using the representative partitions discovered by our KDE-DPC algorithm. Since our main focus is on the broader CBI methodology, and because these tools are thoroughly detailed by Balocchi and Wade, (2025), we refer the interested reader to their thoughtful discussion.

### v2

@@ 4 Credible region
with $\mathbb{P}_{\boldsymbol{\theta}^{2}\overset{\textnormal{iid}}{\sim}\Pi}$-probability at least $1-\delta$, implying $O(N^{-1/2})$ concentration around the desired coverage level $1-\alpha$. For a high-probability guarantee that the coverage merely exceeds $1-\alpha$, one may instead use the results in the references cited above.
4.1 Connections to metric balls as credible regions
Given a discrepancy $\mathcal{D}$ and a point estimate $\hat{\theta}\in\Theta$, a natural alternative credible set is the $\mathcal{D}$-ball around $\hat{\theta}$ with radius chosen to achieve (empirical) $1-\alpha$ posterior coverage. For instance, in Bayesian clustering with the VI metric, this is the standard approach popularized by wade2018clustering. Intriguingly, such credible balls can be recovered within our CBI framework by simply choosing a different scoring rule. Let $\hat{\theta}$ be any estimator computed from the training sample $\boldsymbol{\theta}^{1}$ (e.g., the $\mathcal{D}$-KDE maximizer within $\boldsymbol{\theta}^{1}$). For each calibration draw, define
[…]
Standard conformal prediction results (vovk2005algorithmic, angelopoulos2023conformal) then yield the $(1-\alpha)$ conformal credible region
$\{\theta\in\Theta:\mathcal{D}(\theta,\hat{\theta})\leq\hat{q}_{1-\alpha}\},$
where $\hat{q}_{1-\alpha}$ is the empirical $n^{-1}\lceil(n+1)(1-\alpha)\rceil$-quantile of $\{\mathcal{D}(\theta,\hat{\theta}):\theta\in\boldsymbol{\theta}^{2}\}$. Thus, credible balls emerge as a special case of CBI, but now equipped with the exact, assumption-free posterior coverage guarantees of Theorem 1, rather than simply relying on Monte Carlo validity heuristics.
While this reinterpretation is conceptually useful, we argue that the $\mathcal{D}$-KDE procedure remains preferable in most applications, unless there is strong reason to believe that the posterior is essentially unimodal around $\hat{\theta}$. The intuition behind this is that, on the one hand, our KDE-based conformity score aggregates individual distances to the training samples, so a parameter value receives a relatively high score whenever it is close to any of the modes. This allows the resulting conformal set to track the different high-density regions of the posterior: it can reach several separated modes without having to include the low-density “valleys” between them. On the other hand, a ball centered at a single estimate $\hat{\theta}$ (say, the main mode) must “expand in all directions” until it captures the desired fraction of calibration draws. If the posterior is multimodal, this may have two undesirable consequences. First, to reach the secondary modes it must expand through the low-density regions separating them from $\hat{\theta}$. Second, large portions of the parameter space that are far from all modes may be included, simply as an artifact of the fact that the ball radius must grow uniformly in all directions away from $\hat{\theta}$ until coverage is reached. In either case, the resulting credible set will be unnecessarily large and include areas of the parameter space that are not representative of the posterior support.
Figure 2 illustrates these points with a one-dimensional continuous distribution. The credible sets centered at either the sample mean or the KDE mode become unnecessarily large: both include the low-density region between the modes, and the set centered at the KDE mode extends far to the left of both modes. The KDE-based conformal set, by contrast, adapts perfectly to the multimodal distribution and excludes the unrepresentative areas. Example 2 below shows that these phenomena persist in the substantially more complex space of data partitions.
Figure 2: One-dimensional illustration comparing the KDE-based credible set with Euclidean balls centered at point estimates (all with 90% coverage). The KDE-based set adapts to multimodality, whereas ball-type sets include large low-density regions.
4.2 Theoretical justification of iid assumption
In this subsection, we provide theoretical justification for the iid assumption on the calibration scores. Before doing so, we observe that even if this assumption fails, one can still attach probabilistic guarantees to the coverage achieved by $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$. Indeed, because the ordering of the calibration scores plays no role in constructing the credible region, we may without loss of generality randomly shuffle them and regard the resulting sequence as exchangeable. Since exchangeability is precisely the condition required for the conformal guarantees to hold, the marginal coverage guarantee in Equation (4) continues to apply, with the iid assumption replaced by the corresponding exchangeable law for the calibration and new samples. The interpretation of the guarantee as a statement about the expected posterior coverage changes accordingly, as does the null hypothesis under which the quantity in Equation 2 is formally a $p$-value. [^8] The high-probability guarantee in Equation 4 can also be adapted, although doing so requires additional care. Nevertheless, to maintain a cleaner and more interpretable connection between the theoretical guarantees and the outputs of our procedure, it is valuable to establish conditions under which the calibration sample may indeed be viewed as iid even when it is generated from MCMC, which we tackle next.
@@ 4.2 Theoretical justification of iid assumption
Recall that we collect N:=TSN:=T-S calibration samples. For simplicity, re-index them as $\theta_{0},\theta_{M},\ldots,\theta_{M(N-1)}$, where $M$ acts as a thinning parameter determining which samples are included in the calibration set (that is, every $M$th element from the original sample is retained). While the iid assumption is crucial for obtaining a valid credible region under conformal prediction principles, in practice we often have access not to iid samples from $\Pi$, but to Markov-dependent samples produced by an MCMC algorithm. However, such a sampler is typically designed to be $\Pi$-mixing, meaning that $\lim_{t\to\infty}\sup_{\theta_{0}\in\Theta}\|\mathcal{L}_{t}(\cdot\mid\theta_{0})-\Pi\|_{TV}=0,$ where $\mathcal{L}_{t}(\cdot\mid\theta_{0})$ denotes the distribution of $\theta_{t}$ conditional on the chain starting at $\theta_{0}$, and $\|\cdot\|_{TV}$ denotes the total variation (TV) distance between probability measures. Recall that, for two probability measures $\mu$ and $\nu$ on a sample space $\Omega$, the TV distance is defined as
@@
Under stronger assumptions, one can sometimes establish quantitative convergence, i.e., that the chain is $\Pi$-mixing at rate $\varepsilon_{t}$, meaning that $\sup_{\theta_{0}\in\Theta}\|\mathcal{L}_{t}(\cdot\mid\theta_{0})-\Pi\|_{TV}\leq\varepsilon_{t},$ for some decreasing sequence $(\varepsilon_{t})_{t\in\mathbb{N}}$; see levin2017markov for a standard and comprehensive treatment.
The following result shows that, as long as the chain is mixing, the joint law of the calibration scores $(s_{0},s_{M},\ldots,s_{M(N-1)})$ computed from a suitably thinned calibration sample $(\theta_{0},\theta_{M},\ldots,\theta_{M(N-1)}),$ approaches that of a posterior iid vector in TV distance (as $M\to\infty$).
@@
In practice, Proposition 2 ensures that, even if the sampled parameters are Markov dependent, [^9] Notice that, in our random partition application, the MCMC algorithm operates on a higher-dimensional space than $\Theta$ (i.e., including mixture parameters in addition to the data partition). The resulting chain is then Markov on this larger space, but not necessarily when projected on $\Theta$. This does not affect Proposition 2, since the scores can be viewed as functions from the higher-dimensional state space to the real line. For simplicity, however, we state the result assuming the chain is defined directly on $\Theta$. appropriate thinning allows one to treat them as effectively iid. In fact, if the chain is $\Pi$-mixing at rate $\varepsilon_{t}$, one can even choose the spacing $M$ to achieve a desired coverage level in light of Proposition 1: by the definition of the TV distance in terms of bounded test functions, Proposition 2 implies that selecting $M$ large enough ensures an expected value of $\Pi(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2}))$ close to $1-\alpha$ (as $\boldsymbol{\theta}^{2}\mapsto\Pi(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2}))$ is bounded by 1). Similarly, using the set-based definition of TV distance, one can tune $M$ to obtain an approximate high-probability coverage statement.
We conclude this discussion by noting that the linear dependence on $N$ of the bound in Equation (7) arises from a rather crude argument that treats the calibration samples as ordered and thus Markov dependent. In practice, however, we have already shown that the scores may be seen as exchangeable for the purposes of building $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$. From this viewpoint, a larger $N$ can even improve the TV bound, since shuffling more samples tends to reduce dependence more effectively. We do not pursue this refinement further, as it adds little methodological insight beyond the main point that, after appropriate thinning and burn-in, the calibration samples can be safely regarded as iid draws from the target posterior distribution.
@@
Consider the simulated data and MCMC output analyzed in Example 1. Based on the $\mathcal{D}_{VI}$–KDE calibration scores, we construct the credible set $\mathcal{C}_{0.9}(\boldsymbol{\theta}^{2})$. To assess its properties, we test whether the following partitions belong to it: (i) the true data partition shown in Figure 1(a); (ii) the same (collapsed) partition with the two leftmost clusters merged; (iii) 1,000 randomly generated partitions, each drawn independently by first sampling $K$ from the empirical posterior distribution of the number of clusters (cf. Figure 1(b)), and then assigning each observation to one of $K$ clusters uniformly at random.
@@
We find that both partitions in (i) and (ii) are contained in the set. Given the point estimate shown in Figure 1(c), the inclusion of the collapsed partition is unsurprising, as it closely resembles that estimate. More interestingly, the true partition itself is also included in the set, even though point estimation alone would not have led us to expect this, since the true and collapsed partitions are quite different. This indicates the presence of posterior uncertainty, likely in the form of multimodality, which we will examine further in the next section. Moreover, none of the 1,000 randomly generated partitions fall within the credible region. This supports the view that the credible set does not achieve high coverage simply by encompassing many arbitrary partitions (e.g., those consistent with the marginal posterior on $K$ but otherwise unstructured), but rather by concentrating on representative regions of the posterior.
As a further demonstration of the reasonable size of our credible set, and as an illustration of the advantages of our $\mathcal{D}_{VI}$-KDE region over $\mathcal{D}_{VI}$-balls (centered at the $\mathcal{D}_{VI}$-KDE point estimate, essentially as proposed by wade2018clustering; cf. Subsection 4.1), we examine how both procedures behave on the two partitions displayed in Figure 3. The partition on the left can be reached by “moving away” from the collapsed partition in the “opposite direction” compared to the true partition (recall that both the true and collapsed partitions lie in our credible set, and in Section 5 below we will show that they practically coincide with the two posterior modes), as it groups into the leftmost cluster of the collapsed partition several observations that clearly belong to the rightmost one. The partition on the right of Figure 3, by contrast, lies “between” the two modes: it shrinks one of the true leftmost clusters of the true partition by reassigning some of its members to the other left cluster, thereby making the partition closer to the collapsed one.
In both scenarios, our $\mathcal{D}_{VI}$-KDE conformal $p$-value falls below the $0.1$ inclusion threshold, whereas the $p$-value associated with the credible ball remains above the threshold. This indicates the presence of low-density regions both between and away from the two modes, and while our KDE-based procedure correctly downweights these regions by focusing on the effective support of the posterior, the more crude credible ball unnecessarily includes the two unrepresentative partitions coming from such regions. This behavior is fully consistent with the one-dimensional intuition illustrated earlier in Figure 2.
Figure 3: Two partitions tested according to the $\mathcal{D}_{VI}$-KDE and and $\mathcal{D}_{VI}$-ball conformal procedures. The partition on the left moves away from the collapsed partition in the opposite direction with respect to the true partition, while the partition on the right lies between the true and collapsed partitions.
5 Multimodality analysis
@@
Having defined our approach to point estimation and credible region construction, we now turn to the last piece of the CBI program, that is to address multimodality in the posterior distribution over $\Theta$. Importantly, because the parameter spaces embodied by $\Theta$ are generally non-Euclidean, we require a notion of multimodality that is operationally meaningful within these structures. In particular, we have defined a pseudo-density function, based on the training data, that can be evaluated at any $\theta\in\Theta$, and we can compute a distance $\mathcal{D}(\theta,\theta^{\prime})$ between any two parameters $\theta,\theta^{\prime}\in\Theta$. These two ingredients are sufficient to define a concept of multimodality grounded in the idea of density-based clustering. The latter provides a general framework, encompassing many algorithmic instantiations (ester1996density, kriegel2011density, rodriguez2014clustering), based on the following principle: given a notion of density [^10] The term “density” is somewhat ambiguous in this context, as it may refer either to a notion of local concentration of mass (or sampled points) or, relatedly when applicable, to the probabilistic concept of a density as a Radon–Nikodym derivative with respect to a dominating measure. In our setting, the former interpretation is the most appropriate. on $\Theta$ and a measure of discrepancy between its elements, one partitions the data into clusters such that (a) each cluster lies in a high-density region whose elements are mutually close, and (b) points from different clusters are far apart. A distribution is then said to be multimodal if random samples drawn from it fall into more than one cluster.
@@
Among the many algorithms following this idea, we take inspiration from the Density Peak Clustering (DPC) procedure by rodriguez2014clustering, which is simple, computationally efficient, and well suited to our setting. In particular, since the construction of the credible region already requires the NN $\mathcal{D}$-KDE calibration scores to build the credible set, this algorithm can be applied with minimal additional computation. The only remaining step is to compute the pairwise distances $\mathcal{D}(\theta,\theta^{\prime})$ between all distinct calibration samples $\theta,\theta^{\prime}$; this has $O(N^{2})$ cost, but is parallelizable across each distance computation. Then our procedure, which we term KDE-DPC, proceeds as follows:
@@
    For each calibration parameter $\theta\in\boldsymbol{\theta}^{2}$, compute
    $\delta(\theta):=\min_{\theta^{\prime}\in\boldsymbol{\theta}^{2}\,:\,s(\theta^{\prime};\boldsymbol{\theta}^{1})>s(\theta;\boldsymbol{\theta}^{1})}\mathcal{D}(\theta,\theta^{\prime}),$
    that is, the minimum distance from $\theta$ to any other parameter with higher density. For the parameter(s) with highest density (corresponding to our point estimate $\theta_{\star}$), set by default
    $\delta(\theta_{\star}):=\max_{\theta^{\prime}\in\boldsymbol{\theta}^{2}}\mathcal{D}(\theta_{\star},\theta^{\prime}).$
@@
    Identify the potential cluster centers as those calibration parameters exhibiting simultaneously high density $s(\theta;\boldsymbol{\theta}^{1})$ and large separation $\delta(\theta)$.
@@
We note that, if one were to follow rodriguez2014clustering exactly, the density for each calibration sample $\theta$ would instead be estimated as the proportion of samples within a fixed-radius $\mathcal{D}$-neighborhood of $\theta$. While feasible, this would waste computation, since the KDE scores have already been computed, and it would ignore the information contributed by the training samples $\boldsymbol{\theta}^{1}$. Nevertheless, the overall approach is retained, as the $\mathcal{D}$-KDE remains distance-based, with the exponential kernel acting as a smoother version of a strict neighborhood.
The KDE-DPC procedure exhibits several appealing features. First, as illustrated in Example 3, the selection of representative parameter values corresponding to posterior modes reduces to a simple visual task: inspecting a decision graph where, for each calibration sample $\theta$, the quantities $\delta(\theta)$ (on the vertical axis) and $s(\theta;\boldsymbol{\theta}^{1})$ (on the horizontal axis) are plotted. The posterior modes are then identified as those samples that stand out as having abnormally large values of both quantities—that is, points lying in the north-east corner of the decision graph. This approach has the practical advantage of not requiring a pre-specified the number of modes, which is instead discovered together with the modes themselves. Second, the posterior weight associated with each mode can be naturally estimated as the proportion of calibration samples assigned to its cluster. Third, the method provides a straightforward way to classify calibration samples as inliers or outliers: samples lying in the north-west corner of the decision graph can be interpreted as outliers, since they are far from higher-density samples yet have low density themselves, while those located in the middle and south-east regions are inliers, as they exhibit moderate to high density and are close to representative modes. These considerations significantly enhance one’s understanding of the uncertainty encoded by the posterior distribution.
@@
We continue our analysis of the MCMC samples of random partitions introduced in Examples 1 and 2. Figure 4(a) displays the KDE-DPC decision graph, plotting for each calibration partition its separation $\delta$ and density $s(\cdot;\boldsymbol{\theta}^{1})$ values, with points colored by the number of clusters in the corresponding partition. [^11] A large number of two-cluster partitions in the calibration set coincide exactly with the partition corresponding to the north-east-most point, which explains the limited number of visible blue points. Two points (highlighted in red) clearly stand out in the north-east corner of the graph, identifying the representative modes. The corresponding partitions, shown in Figure 4(b) (ordered by density from left to right), indicate that the posterior concentrates substantial mass around two distinct configurations: one with two clusters, obtained by merging the two leftmost clusters of the true data partition (coinciding by default with $\theta_{\star}$ as found in Example 1), and another nearly identical to the true partition itself. [^12] See the Supplementary Material for an additional experiment using the same simulated data and a similar prior, where the posterior over partitions is instead found to be unimodal.
Interestingly, coloring by the number of clusters reveals that, while the posterior assigns non-negligible mass to partitions with four clusters, none of these emerge as representative modes. Instead, they appear either as inliers close to one of the two main modes (the yellow points at the center of the decision graph) or as outliers lying in low-density regions far from the modes (the yellow points in the north-west area of the decision graph). In this sense, the multimodality analysis suggests that, although four-cluster configurations receive some posterior support under the PY mixture model used to fit the data, they are likely spurious from the standpoint of a parsimonious description of the uncertainty encoded in the posterior distribution.
@@
To conclude, we note that the output of our multimodality analysis based on the KDE-DPC procedure is analogous to that of balocchi2025understanding, who summarize the posterior over partitions using a Wasserstein-distance-minimizing mixture of point masses at representative partitions. In both approaches, the result is a set of posterior modes capturing meaningfully separated regions of posterior concentration in the partition space. A few advantages of our approach are worth highlighting. First, it only requires density scores (already computed for the credible region) and pairwise VI distances between calibration samples (also required by balocchi2025understanding), both of which can be parallelized to achieve substantial computational gains. In contrast, the $k$-medoids-like algorithm of balocchi2025understanding is inherently sequential. Second, their method requires specifying the number of modes in advance. To select a good value of this important hyperparameter, the authors propose running the algorithm for different choices and picking the one with the largest improvement in posterior approximation. Clearly, this approach may be computationally cumbersome and prone to overfitting to the available MCMC samples. Our KDE-DPC procedure, by contrast, automatically reveals the number of modes through visual inspection of the decision graph. Finally, balocchi2025understanding propose several insightful techniques to characterize posterior uncertainty once the modes are identified, such as plotting the meet between any two representative partitions or examining the VI contribution of each data point. All of these analyses can naturally be performed using the representative partitions discovered by our KDE-DPC algorithm. Since our main focus is on the broader CBI methodology, and because these tools are thoroughly detailed by balocchi2025understanding, we refer the interested reader to their thoughtful discussion.